Solve for x:
a) (2x-1)(^1/2)=sqrt3
b) 8(^2x)+8(^x) -12=0

Answer by edjones(8007)

a)
(2x-1)(^1/2)=sqrt(3)
sqrt(2x-1)=sqrt(3)
2x-1=3 square each side.
2x=4
x=2
.
b)
8^(2x)+8^x-12=0
(8^x)^2+8^x-12=0
(8^x+4)(8^x-3) Factor
8^x=-4 (extraneous), 8^x=3
log(8)3=.528321
8^x=8^.528321
x=.528321